OpenStudy (anonymous):

Can someone help me solve this? -4(2-y)+5x=6(x-y)

OpenStudy (shamil98):

Use the distributive property. -4(2-y) + 5x = 6(x-y) -8 + 4y + 5x = 6x -6y 4y + 5x = 6x - 6y - 8 10y - x = -8

OpenStudy (shamil98):

for x -x = -10y - 8 x = 10y + 8 for y 10y = x -8 y = x/10 - 4/5

OpenStudy (anonymous):

But I thought you cant have x/10 so how would I solve it on from 10y-x=-8

OpenStudy (anonymous):

@Shrina why do you think we can't have x/10 ??

OpenStudy (anonymous):

he has explained very well

OpenStudy (anonymous):

because when you plug it in the equation, it doesn't make sense

OpenStudy (anonymous):

in which equation you are plugging the value of x??